# Topological signatures of collective dynamics and turbulent-like energy cascades in active granular matter

### v1

@@ Abstract
Active matter refers to a broad class of non-equilibrium systems where energy is continuously injected at the level of individual “particles.” These systems exhibit emergent collective behaviors that have no direct thermal-equilibrium counterpart. Their scale ranges from micrometer-sized swarms of bacteria to meter-scale human crowds. In recent years, the role of topology and self-propelled topological defects in active systems has garnered significant attention, particularly in polar and nematic active matter. Building on these ideas, we investigate emergent collective dynamics in apolar active granular fluids. Using granular vibrators as a model experimental system of apolar active Brownian particles in a dry environment, we uncover a distinctive three-stage time evolution arising from the intricate interplay between activity and inelastic interactions. By analyzing the statistics, spatial correlations, and dynamics of vortex-like topological defects in the displacement vector field, we demonstrate their ability to describe and predict this intrinsic collective motion. Furthermore, we show that topological defects play a crucial role in the development of a turbulent-like inverse energy cascade, where kinetic energy transfers across different length scales over time. As the system evolves, the power scaling of the energy transfer increases with the duration of observation. Our findings demonstrate how topological concepts can be applied to predict macroscopic collective phenomena in apolar active matter. This establishes a direct link between microscopic topological dynamics and large-scale behaviors in active granular fluids.
@@ Introduction
More recently, activity has been introduced into quasi-two-dimensional granular systems by placing a layer of granular particles on a vibrating surface, effectively supplying energy to individual particles and allowing them to behave as self-propelling units. A wide range of intriguing phenomena has been observed in these systems, including large-scale collective motion in vibrated polar disks [34] and tapered rods [35], giant number density fluctuations in swarming granular rods [36], dispersionless vibrational modes [37] and gapped shear waves [38] in bidisperse nonpolar Brownian vibrators, as well as high-energy velocity tails in granular spinners [39] and monodisperse nonpolar Brownian vibrators [40]. Additionally, flocking behavior has been observed in monodisperse nonpolar Brownian vibrators [41]. These systems, commonly referred to as active granular matter, form the central focus of our work.
@@
Figure 1: Experimental setup and underlying physics. (a) Images of the large and small active Brownian particles used in our experiment. Left: Photos showing the top and bottom views of both particle types. Right: A schematic diagram illustrating the particle structure with relevant dimensions. (b) An image of the electromagnetic shaker, with a layer of bi-disperse particles placed on an aluminum plate mounted on the shaker. (c) When subjected to vertical vibrations, particles gain kinetic energy individually. This activity induces random collisions, primarily between one of their legs and the bottom plate, leading to isotropic random walk motion in the horizontal 2D plane. (d) Displacement vector field of bi-disperse particles at a low packing fraction ($\phi=0.309$), where activity dominates leading to uncorrelated gas-like dynamics. (e) A schematic diagram illustrating how collective dynamics emerge from particle-particle inelastic interactions that become dominant at high packing fractions. (f) Displacement field of bi-disperse particles at a high packing fraction ($\phi=0.822$), exhibiting collective vortex-like structures driven by interparticle interactions.
@@
Historically, topology has played a central role in studying polar and nematic active systems, where anisotropic shapes or directional self-propulsion define preferred orientations. Familiar concepts from liquid crystals [70] apply directly in these contexts. Collective motion, typically associated with self-aligning interactions—as exemplified by the Vicsek model [71]—was long thought to be restricted to such systems. However, recent theoretical work [72], later confirmed experimentally [41], demonstrated that flocking and large-scale motion can also arise in active granular systems without self-aligning interactions. In these systems, effective attractive forces emerge from inelastic collisions, driving collective behavior.
@@
This work aims to extend the role of topology in nonpolar active granular systems without self-aligning interactions, proving a direct connection between topological defect dynamics, large-scale collective motion, and turbulent-like structures.
Active granular fluids
@@
In Figure 1, we present the most salient features of our experimental setup and the essential ingredients to understand the underlying physics.
@@
The microscopic constituents in our experimental platform are active Brownian particles that were 3D printed using a resin material. Each particle consists of a disk-shaped cap with $12$ staggered legs attached underneath, as shown in Fig. 1(a). We utilize bidisperse particles with radii of $R_{l}=11.2$ mm for large particles and $R_{s}=8$ mm for small particles. The other geometric characteristics are the same for large and small particles, namely disk thickness ($h_{1}=3$ mm), leg height ($h_{2}=3$ mm), and placement of $12$ staggered legs along a concentric circle of radius $r=5.6$ mm. The mixing ratio between large and small particles is $1:2$. Given their size ratio ($R_{l}:R_{s}=1.4:1$), the total area occupied by large and small particles remains equal, maintaining a $1:1$ area ratio in the whole system.
@@
A single layer of particles is then placed on an aluminum plate within a flower-shaped region mounted on top of an electromagnetic shaker, see Figure 1(b). The electromagnetic shaker generates a vertical sinusoidal vibration at $100$ Hz, driving the entire system at the individual particle level with a maximum acceleration of $2.5g$, where $g$ is the gravitational acceleration. A Basler CCD camera (not shown in the image) is positioned above the experimental platform and continuously records particle configurations for an hour at a rate of $40$ frames per second. Figure 1(d, f) provides two images of the particles captured by the camera at different packing fractions.
@@
During an experimental run, each particle individually receives kinetic energy from the shaker’s vertical vibration and dissipates it through collisions with other particles and with the plate mounted on the shaker. Once the energy injection and the dissipation balance on average, a non-equilibrium steady state is reached. The system’s complex dynamics and collective motions arise from the interplay between individual particle activity and interparticle inelastic interactions, as illustrated in the schematic representations in Fig.1(c) and (e).
@@
The competition between these effects gives rise to two distinct dynamical behaviors, observed at low and high packing fractions (panels (d) and (f) in the same Figure). The packing fraction is defined as the ratio of the area occupied by the particles to the total area of the system. At low packing fractions, activity is dominant. Fig.1(d) shows the random displacement field superimposed on the individual particles within this regime, at a low packing fraction $\phi=0.309$. Due to the unique leg design, a single particle gains kinetic energy through stochastic collisions between its legs and the bottom plate, resulting in random walk motion along the two-dimensional horizontal plane. This mechanism acts independently on each particle, giving rise to random gas-like motion as illustrated in Fig.1(c).
@@
In this dilute limit, the dynamics are uncorrelated, and a single particle’s translational and rotational motions follow Gaussian distributions with zero mean, as demonstrated in previous studies [40]. Thus, these particles can be considered a granular realization of active Brownian particles [73]. On the other hand, as shown in Fig.1(e)), inelastic interactions tend to correlate the dynamics of the individual particles, leading to collective motion involving clusters of neighbor particles. This effect becomes more pronounced by increasing the packing fraction. In the sufficiently dense regime, particle-particle interactions dominate over individual particle activity, giving rise to collective motion. At a high packing fraction ($\phi=0.822$), large collective structures, with vortex-like form, appear in the particle displacement field, as illustrated in Fig.1(f).
@@
Figure 2: Topological defects characterize the self-organization and the onset of collective motion. (a) Displacement field (black arrows with thickness representing magnitude) for different observation times, $\tau=0.05\,s$, $10\,s$ and $500\,s$. Red and blue symbols refer to vortices and anti-vortices with ±$\pm 1$ winding number $q$. The color map represents the local value of the polarization $P_{c}$. As a reference, the white bar shows the length-scale $2\,D_{s}$, with $D_{s}$ the diameter of the small particles. (b) Snapshots of two consecutive displacement frames highlighting the annihilation of a $+1$ vortex (red) with a $-1$ anti-vortex (blue), conserving the total topological charge $q_{\text{tot}}$ (00 in the top panel and $+1$ in the bottom one). (c) Same representation as in panel (a) but with the color map indicating the local value of the vorticity intensity $\Omega_{l}$.
@@
To facilitate the computation of relevant physical quantities characterizing self-organization and collective motion, we first discretize the raw particle displacement fields (such as those shown in Fig. 1(d,f)) by mapping them onto square lattice of size $1.84\,D_{s}$. More precisely, a raw displacement field is defined as $\textbf{u}_{i}\equiv\textbf{r}_{i}(t_{0}+\tau)-\textbf{r}_{i}(t_{0})$, where $\textbf{r}_{i}(t)$ denotes the position of particle $i$ at time $t$. Here, $i$ indexes the particles, $t_{0}$ represents a specific reference time, and $\tau$ is the duration of observation time.
@@
By construction, the parameter $P_{c}$ goes from 00, when particle displacements are random and uncorrelated, to $1$, when all displacements point locally toward the same direction. In a sense, $P_{c}$ quantifies the degree of local flocking in the particle dynamics. If $n$ is taken to be the total number of particle $n=N$, then $P_{c}$ describes the onset of global flocking as usually employed in polar active systems, e.g., Vicsek model [71].
@@
[…]
where $\theta$ is the direction (phase) of the local displacement vector concerning the $x$-axes and $\mathcal{C}$ a closed loop. $q=\pm 1$ corresponds respectively to a vortex/anti-vortex where the phase of the displacement vector is not well-defined, and the vector itself is singular. $q$ is a topological conserved charge, insensitive to continuous transformations of the displacement field. It follows that, in a freely evolving system, a vortex can disappear only through annihilation with a corresponding anti-vortex, conserving the total topological charge of the system.
@@
Figure 3: Structural relaxation towards the onset of collective motion is mediated by topological defects. (a) Snapshot of $D_{\text{min}}^{2}$ field (gray color map) with topological defects (red and blue symbols) at the observation time of $\tau=500$ s. The characterization of local collective motion for the same $\tau$ is provided by the third panel in Fig. 2(a). (b) Pair correlation functions between areas of strong structural relaxation, soft spots (SS) (top $5\%$ $D_{\text{min}}^{2}$ particles), and positive topological defects at time interval $\tau$. (c) The same correlation functions for negative topological defects. Each curve has been shifted up by two units on the vertical axis for clearer visualization.
@@
In Fig.2, we present three snapshots of the particle displacements on the square lattice for three different observation times: $\tau=0.05\,s$, $\tau=10\,s$, and $\tau=500\,s$. The local displacement vectors are visualized with black arrows whose thicknesses represent their amplitudes. These different timescales are taken as benchmark values to characterize our system’s short-, intermediate-, and long-term dynamics.
@@
Looking at the spatial structure of the displacements, it is clear that at short time scales ($\tau=0.05$ s in Fig.2), the motion of the particles is uncorrelated, random and independent, as in a gas-like environment. Kinematically, this is the consequence of the activity injecting kinetic energy at the individual particle level and favoring a two-dimensional (2D) Brownian motion similar to the effect of temperature in thermal systems. Observing the dynamics on longer time scales makes the emergence of collective motion and large-scale cooperative structures evident. As we argue below, this emergence can be ascribed to the inelastic interparticle collisions and rationalized by topological concepts applied to the displacement vector field.
@@
In Fig.2(a), we show the value of the local displacement polarization $P_{c}$ using a color map ranging from blue ($P_{c}=0$) to yellow ($P_{c}=1$). Thus, yellow regions represent clusters with strong collective motion, where the particle displacements are aligned as flocks of birds. As the observation time increases, from the left to right panels, the size of the cooperative (yellow) regions grows, indicating the emergence of larger-scale collective motion. To provide a microscopic explanation of this behavior, we present the location of the topological defects (TDs) in the same panels, using red symbols for vortices and blue symbols for anti-vortices. We notice that at short times ($\tau=0.05$ s), the topological defects are located in random positions, forming a gas of defects. Moreover, TDs tend to be located in regions with low polarization, indicating that TDs are microscopic objects that disrupt the onset of collective motion by randomizing the direction of the particle displacements. More details on the spatical correlation between TDs as a function of the observation time $\tau$ can be found in Appendix B.
@@
Importantly, we observe that the onset of collective motion, characterized by the percolation of clusters with large $P_{c}$, is mediated by the annihilation of topological defects that ‘live’ at the interfaces between these collective clusters (blue regions). This annihilation occurs between pairs of defects with opposite charges, preserving the total winding number $q_{\text{tot}}$. Two examples of this process are shown with enlarged views in Fig.2(b). It is also evident and discussed in more detail below that the number of TDs decreases with the observation time, favoring the establishment of large-scale collective clusters.
@@
To provide a complementary view of these dynamics, in Fig.2(c), we present the same particle displacements and TDs but, this time, superimposed onto the amplitude of the local displacement curl $\Omega_{l}$. As expected, the local curl color map (Fig.2(b)) and the local polarization color map (Fig.2(a)) are roughly complementary. Moreover, the local curl $\Omega_{l}$ is significant around the TDs, that can therefore be identified as the microscopic objects disrupting the irrotational nature of the displacement field. In this complementary view, the onset of collective motion is driven by the percolation of areas with low local displacement curl, which again is driven by the annihilation of defect pairs (Fig.2(c)).
@@
Figure 4: Three stage non-equilibrium dynamics. (a): Topological signatures. Defect statistics as a function of the observation time $\tau$. $N_{\text{defects}}$ is the total number of TDs. The vertical dashed lines indicate the two time scales $t_{1}=3\,s$, $t_{2}=60\,s$. (b): Particle-level dynamics. Mean square displacement (MSD) in units of the square of the small particle diameter $d_{s}^{2}$. (c): Collective motion. Average curl of the system $\Omega_{0}$. The three red arrows at each panel correspond to the selected times shown in Fig.2.

### v2

# Topological signatures of collective dynamics and turbulent-like energy cascades in apolar active granular matter

@@ Abstract
Active matter refers to a broad class of non-equilibrium systems where energy is continuously injected at the level of individual “particles.” These systems exhibit emergent collective behaviors that have no direct thermal-equilibrium counterpart. Their scale ranges from micrometer-sized swarms of bacteria to meter-scale human crowds. In recent years, the role of topology and self-propelled topological defects in active systems has garnered significant attention, particularly in polar and nematic active matter. Building on these ideas, we investigate emergent collective dynamics in apolar active granular fluids. Using isotropic granular vibrators as a model experimental system of apolar active Ornstein-Uhlenbeck particles in a dry environment, we uncover a distinctive three-stage time evolution arising from the intricate interplay between activity and inelastic interactions. By analyzing the statistics, spatial correlations, and dynamics of vortex-like topological defects in the displacement vector field, we demonstrate their ability to describe this intrinsic collective motion. Furthermore, associated to these topological defects, we reveal the onset of a turbulent-like inverse energy cascade, where kinetic energy transfers across different length scales over time. As the system evolves, the power scaling of the energy transfer increases with the duration of observation. Our findings show that topological concepts can be extended to the nonequilibrium dynamics of apolar active matter, revealing a direct link between microscopic topological processes and emergent large-scale behaviors in active granular fluids that lack both a well-defined direction of motion and an intrinsic axis of orientation at the particle scale.
@@ Introduction
More recently, activity has been introduced into quasi-two-dimensional (2D) granular systems by placing a layer of granular particles on a vibrating surface, effectively supplying energy to individual particles and allowing them to behave as self-propelling units. A wide range of intriguing phenomena has been observed in these systems, including large-scale collective motion in vibrated polar disks [34] and tapered rods [35], giant number density fluctuations in
swarming granular rods [36], dispersionless vibrational modes [37] and gapped shear waves [38] in bidisperse apolar Brownian vibrators, as well as high-energy velocity tails in granular spinners [39] and monodisperse apolar Brownian vibrators [40]. Additionally, flocking behavior has been observed in monodisperse apolar Brownian vibrators [41]. These systems, commonly referred to as active granular matter, form the central focus of our work.
Figure 1: Experimental setup and underlying physics. (a) Images of the large and small active isotropic granular particles used in our experiment. Left: Photos showing the top and bottom views of both particle types. Right: A schematic diagram illustrating the particle structure with relevant dimensions. (b) An image of the electromagnetic shaker, with a layer of bi-disperse particles placed on an aluminum plate mounted on the shaker. (c) When subjected to vertical vibrations, particles gain kinetic energy individually. This activity induces random collisions, primarily between one of their legs and the bottom plate, leading to isotropic random walk motion in the horizontal 2D plane. (d) Displacement vector field of bi-disperse particles at a low packing fraction ($\phi=0.309$), obtained at $t_{0}=0.025\,s$ and $\tau=500\,s$, with their lengths rescaled by a factor of 0.1 for clarity, where activity dominates leading to uncorrelated gas-like dynamics. (e) A schematic diagram illustrating how collective dynamics emerge from particle-particle inelastic interactions that become dominant at high packing fractions. (f) Displacement field of bi-disperse particles at a high packing fraction ($\phi=0.822$), obtained at $t_{0}=0.025\,s$ and $\tau=500\,s$, with their lengths rescaled by a factor of 5 for clarity, exhibiting collective vortex-like structures driven by interparticle interactions. Here, $t_{0}$ denotes a specific initial reference time, and $\tau$ represents the duration of the observation period $t_{0},t_{0}+\tau]$.
@@
Historically, topology has played a central role in studying polar and nematic active systems, where anisotropic shapes or directional self-propulsion define preferred orientations. Familiar concepts from liquid crystals [70] apply directly in these contexts. Collective motion, typically associated with self-aligning interactions—as exemplified by the Vicsek model [71]—was long thought to be restricted to such systems. However, recent theoretical work [72], later confirmed experimentally [41], demonstrated that flocking and large-scale motion can also arise in active granular systems without self-aligning interactions. In these systems, effective attractive forces emerge from inelastic collisions, driving collective behavior. Still it remains unclear [73], whether these systems exhibit all the characteristic features of flocking.
@@
This work aims to extend the role of topology in apolar isotropic active granular systems without self-aligning interactions, proving a direct connection between topological defect dynamics, large-scale collective motion, and turbulent-like structures.
Active granular fluids and single particle dynamics
@@
In Fig.1, we present the most salient features of our experimental setup and the essential ingredients to understand the underlying physics.
@@
The microscopic constituents in our experimental platform are isotropic granular particles that were 3D printed using a resin material. Each particle consists of a disk-shaped cap with $12$ staggered legs attached underneath, as shown in Fig. 1(a). We utilize bidisperse particles with diameter of $D_{l}=22.4$ mm for large particles and $D_{s}=16$ mm for small particles. The other geometric characteristics are the same for large and small particles, namely disk thickness ($h_{1}=3$ mm), leg height ($h_{2}=3$ mm), and placement of $12$ staggered legs along a concentric circle of radius $r=5.6$ mm. Each leg is inclined inward at an angle of 18.4°, alternating with a deviation of ±38.5° from the particle’s mid-plane. The mixing ratio between large and small particles is set to $1:2$ to suppress crystallization. Given their size ratio ($D_{l}:D_{s}=1.4:1$), the total area occupied by large and small particles remains equal, maintaining a $1:1$ area ratio in the whole system.
@@
A single layer of particles is then placed on an aluminum plate within a flower-shaped region mounted on top of an electromagnetic shaker, see Fig.1(b), designed to suppress collective rotation along the boundary and to facilitate reinjection of particles near the edges back into the system [74]. A detailed discussion of the effect of the boundary shape on the final vortex configuration is provided in the Supplementary Information (SI). The electromagnetic shaker generates a vertical sinusoidal vibration at $100$ Hz, driving the entire system at the individual particle level with a maximum acceleration of $2.5g$, where $g$ is the gravitational acceleration. A Basler CCD camera (not shown in the image) is positioned above the experimental platform and continuously records particle configurations for an hour at a rate of $40$ frames per second. Fig.1(d, f) provides two images of the particles captured by the camera at different packing fractions.
Before each experiment, the vibration platform was carefully leveled to suppress gravitational drift. Levelness was confirmed by achieving a uniform particle distribution at low packing fractions after two hours of continuous vibration, well beyond the characteristic drift time of a few seconds. The system was then vibrated for two hours prior to data collection to establish a steady initial state, ensuring reproducible and well-controlled conditions.
During an experimental run, each particle individually receives kinetic energy from the shaker’s vertical vibration and dissipates it through collisions with other particles and with the plate mounted on the shaker. Particle collisions are inelastic, characterized by a restitution coefficient of $\epsilon\approx 0.39$ (see SI for measurement details). The static friction coefficient between the particles and the plate is $\mu_{s}\approx 0.42$ [41]. The collision between a granular vibrator and the plate acts as the source of the active force. Under external vibration, collisions between each granular vibrator and the plate are effectively collisions between individual legs and the plate. In each event, typically only one leg contacts the substrate. Due to the inclined geometry, every collision imparts a net displacement to the particle in a specific direction, creating a propulsion mechanism that mimics an active force.
@@
Once the energy injection and the dissipation balance on average, a non-equilibrium steady state is reached. The system’s complex dynamics and collective motions arise from the interplay between individual particle activity and interparticle inelastic interactions, as illustrated in the schematic representations in Fig.1(c) and (e).
Figure 2: Single particle dynamics. (a) 2D velocity distribution $P(v_{x},v_{y})$ for large particles. (b) 2D velocity distribution $P(v_{x},v_{y})$ for small particles. (c) Single-particle velocity time correlation $C_{v}(t)$. (d) Single-particle orientation time correlation $C_{n}(t)$. In both panels (c) and (d), the red curve represents large particles while the blue curve represents small particles. The fitting functions are for $C_{v}(t)$ and $C_{n}(t)$ are given by a simple exponential $C_{v}(t)=a\exp(-t/b)$ and (1) respectively. The parameters resulting from the fitting procedure are indicated as insets in panels (c) and (d).
@@
The competition between these effects gives rise to two distinct dynamical behaviors, observed at low and high packing fractions (panels (d) and (f) in the same Figure). The packing fraction is defined as the ratio of the area occupied by the particles to the total area of the system. At low packing fractions, activity is dominant. Fig.1(d) shows the random displacement field superimposed on the individual particles within this regime, at a low packing fraction $\phi=0.309$. Due to the unique leg design, a single particle gains kinetic energy through stochastic collisions between its legs and the bottom plate, resulting in random walk motion along the 2D horizontal plane. This mechanism acts independently on each particle, giving rise to random gas-like motion as illustrated in Fig.1(c).
@@
On the other hand, as shown in Fig.1(e)), inelastic interactions tend to correlate the dynamics of the individual particles, leading to collective motion involving clusters of neighbor particles. This effect becomes more pronounced by increasing the packing fraction. In the sufficiently dense regime, particle-particle interactions dominate over individual particle activity, giving rise to collective motion. At a high packing fraction ($\phi=0.822$), large collective structures, with vortex-like form, appear in the particle displacement field, as illustrated in Fig.1(f).
@@
In order to precisely characterize our active system, we begin with the analysis of single-particle dynamics. As shown in Fig. 2(a)–(b), the 2D velocity distribution is centered at the origin, $(v_{x},v_{y})=(0,0)$, for both large and small particles. This indicates that the system lacks a well-defined intrinsic velocity scale.
Furthermore, Fig. 2(c) shows the velocity time-correlation function $C_{v}(t)$ (see Methods), which decays rapidly and monotonically, well described by an exponential law $C_{v}\propto\exp(-t/b)$ with $b=0.045$ s and $0.042$ s for large and small particles, respectively. This demonstrates that velocity correlations are finite and short-lived, with translational dynamics fully decorrelating beyond $\sim 10^{-1}$ s.
Inspired by recent theoretical work [75], we also analyze the orientation correlation $C_{n}(t)$ (see Methods), shown in Fig.2(d). Following Ref. [75], we fit the data to
$C_{n}(t)=\frac{2D_{\chi}t_{p}}{t_{p}^{2}-\tau_{\chi}^{2}}\left(t_{p}\,e^{-t/t_{p}}-\tau_{\chi}\,e^{-t/\tau_{\chi}}\right),$ (1)
where $D_{\chi}$ is the inertial rotational diffusivity, $\tau_{\chi}$ the rotational memory time, and $t_{p}$ a characteristic timescale coinciding with the persistence time in overdamped systems. From the unit normalization of the orientation vector, $D_{\chi}$ can be obtained as [76]:
$D_{\chi}=\frac{t_{p}+\tau_{\chi}}{2\tau}.$ (2)
Fitting our data yields $t_{p,l}=0.113\,\mathrm{s}$ and $\tau_{\chi,l}=0.329\,\mathrm{s}$ for large particles, and $t_{p,s}=0.159\,\mathrm{s}$ and $\tau_{\chi,s}=0.159\,\mathrm{s}$ for small particles. For small particles, the fit agrees reasonably well with theory, though the tail exhibits deviations: the experimental data show an oscillatory trend absent in the theoretical curve.
Taken together, the results of Fig. 2 demonstrate that our system closely resembles active Ornstein–Uhlenbeck particles (AOUPs): it exhibits a finite persistence time, but, unlike active Brownian particles, lacks a characteristic nonzero velocity scale[75, 76].
Moreover, we analyze the effective Péclet number of single particles, defined as the ratio of the persistence length to the particle size [77]:
$\mathrm{Pe}=\frac{v_{0}\tau_{p}}{D},$ (3)
where $v_{0}$ is the active velocity, $\tau_{p}$ the persistence time, and $D$ the particle diameter. We obtain $\mathrm{Pe}_{l}=0.0282$ for large particles and $\mathrm{Pe}_{s}=0.0358$ for small particles. The persistence time is $\tau_{p}=0.025\,\mathrm{s}$ for both particle types, which is much shorter than the typical observation times employed in our experiments $(\tau=0.05\,\mathrm{s},10\,\mathrm{s},\text{and }500\,\mathrm{s})$ (see SI for detailed measurement). Moreover, it is significantly smaller than the relaxation timescales shown in Fig. S5 in the Supplementary Information.
Figure 3: Topological defects characterize the self-organization and the onset of collective motion at the packing fraction $\mathbf{\phi=0.822}$. (a) Displacement field (black arrows with thickness representing magnitude) for different observation times, $\tau=0.05\,s$, 10s10\,s and 500s500\,s. Red and blue symbols refer to vortices and anti-vortices with $\pm 1$ winding number $q$. The color map represents the local value of the polarization $P_{c}$. As a reference, the white bar shows the length-scale $2\,D_{s}$, with $D_{s}$ the diameter of the small particles. (b) Snapshots of two consecutive displacement frames highlighting the annihilation of a +1+1 vortex (red) with a 1-1 anti-vortex (blue), conserving the local topological charge (0 in the top panel and +1+1 in the bottom one). (c) Same representation as in panel (a) but with the color map indicating the local value of the vorticity intensity $\Omega_{l}$.
@@
In this work, unless stated otherwise, our analysis concerns a system of 798 bidisperse particles at a packing fraction of 0.822. Fig.S4(a) in the Supplementary Information shows the intermediate scattering function at this packing fraction, demonstrating that particle displacements remain minimal at late time without any evident relaxation observed within the experimental time window. In this system, particles are densely packed and the density is approximately uniform, as shown in Fig.1(f), forming the basis for the following analysis.
To facilitate the computation of relevant physical quantities characterizing self-organization and collective motion, we first discretize the raw particle displacement fields (such as those shown in Fig. 1(d,f)) by mapping them onto square lattice of size $1.84\,D_{s}$. The coarse-grained displacement field is obtained by summing the displacements of all particles within each lattice cell. More precisely, a raw displacement field is defined as $\textbf{u}_{i}\equiv\textbf{r}_{i}(t_{0}+\tau)-\textbf{r}_{i}(t_{0})$, where $\textbf{r}_{i}(t)$ denotes the position of particle $i$ at time $t$. Here, $i$ indexes the particles, $t_{0}$ represents a specific reference time, and $\tau$ is the duration of observation time. In our analysis, the time interval $\tau$, over which displacements are measured, can be varied freely from $\tau=0.025$ s up to $\tau\sim 10^{3}$ s. To obtain reliable statistics, we divide the time sequence of each experimental run into several intervals of equal length $\tau$. For each interval $t_{0},t_{0}+\tau]$, we compute the displacement field relative to the reference time $t_{0}$, and then collect ensemble-averaged statistics over at least 500 different $t_{0}$ for each analysis to ensure statistical reliability. The advantage of using the displacement field, rather than the velocity field $\mathbf{v}\equiv\lim_{dt\to 0}d\mathbf{u}/dt$, is to probe the dynamics of the system at all timescales and to investigate the systematic effect of particle-particle interactions, something that is not possible by considering only the $dt\to 0$ regime.
@@
By construction, the parameter $P_{c}$ goes from 0, when particle displacements are random and uncorrelated, to $1$, when all displacements point locally toward the same direction. In a sense, $P_{c}$ quantifies the degree of local flocking in the particle dynamics. If $n$ is taken to be the total number of particle $n=N$, then $P_{c}$ describes the onset of global flocking as usually employed in polar active systems, e.g., Vicsek model [71].
@@
$q\equiv\frac{1}{2\pi}\oint_{\mathcal{C}}d\theta,$ (6)
where $\theta$ is the direction (phase) of the local displacement vector concerning the $x$-axes and $\mathcal{C}$ a closed loop. $q=\pm 1$ corresponds respectively to a vortex/anti-vortex where the phase of the displacement vector is not well-defined, and the vector itself is singular. It should be noted that the defects discussed here are not conventional solid-state defects (such as dislocations or Eshelby quadrupoles), but rather orientational defects, with the displacement field acting as the relevant vectorial order parameter.
For systems with a well-defined order parameter and stiffness, such as the XY model or superfluids, the total topological charge $q$ is conserved. In contrast, in fluid systems lacking rigidity and subject to external forces or activity, the conservation of the total topological charge is no longer guaranteed. Nevertheless, energetic arguments suggest that local conservation of $q$ is still favored, so that defects typically appear or disappear in pairs, annihilating with oppositely charged partners. It is also important to note that the very definition of topological defects requires the existence of a continuous vector field. This assumption inevitably breaks down at low packing fractions, where the particle distribution is too dilute to be well approximated by a continuous field. In what follows, we revisit these points in detail and complement them with a direct analysis of the experimental data.
@@
Figure 4: Structural relaxation towards the onset of collective motion is associated with topological defects. The system packing fraction is fixed at $\phi=0.822$, identical to that in Fig.3. (a) Snapshot of $D_{\text{min}}^{2}$ field (gray color map) with topological defects (red and blue symbols) at the observation time of $\tau=500$ s. The characterization of local collective motion for the same $\tau$ is provided by the third panel in Fig. 3(a). The bar indicates a length scale of $2\,D_{s}$, with $D_{s}$ the diameter of the small particles. (b) Pair correlation functions between areas of strong structural relaxation, soft spots ($S$) (top $5\%$ $D_{\text{min}}^{2}$ particles), and positive topological defects at time interval $\tau$. (c) The same correlation functions for negative topological defects. Each curve has been shifted up by two units on the vertical axis for clearer visualization, and the horizontal axis is scaled by $D_{s}$ in (b) and (c).
@@
In Fig.3, we present three snapshots of the particle displacements on the square lattice for three different observation times: $\tau=0.05\,s$, $\tau=10\,s$, and $\tau=500\,s$. The local displacement vectors are visualized with black arrows whose thicknesses represent their amplitudes. These different timescales are taken as benchmark values to characterize our system’s short-, intermediate-, and long-term dynamics. The snapshots in Fig.3 are taken at the initial time $t_{0}=0.025\,\text{s}$, which corresponds to the temporal resolution limit. The files Movie S1.mp4 and Movie S2.mp4 available in the SI Movies show the dynamics composed of snapshots obtained by varying the initial time $t_{0}$, corresponding to the observation times $\tau=0.05\,\text{s}$ and $\tau=10\,\text{s}$, respectively.
@@
Looking at the spatial structure of the displacements, it is clear that at short time scales ($\tau=0.05$ s in Fig.3), the motion of the particles is uncorrelated, random and independent, as in a gas-like environment. Kinematically, this is the consequence of the activity injecting kinetic energy at the individual particle level and favoring a 2D collisional motion. Observing the dynamics on longer time scales makes the emergence of collective motion and large-scale cooperative structures evident. As we argue below, this emergence can be ascribed to the inelastic interparticle collisions and rationalized by topological concepts applied to the displacement vector field.
@@
In Fig.3(a), we show the value of the local displacement polarization $P_{c}$ using a color map ranging from blue ($P_{c}=0$) to yellow ($P_{c}=1$). Thus, yellow regions represent clusters with strong collective motion, where the particle displacements are aligned as flocks of birds. As the observation time increases, from the left to right panels, the size of the cooperative (yellow) regions grows, indicating the emergence of larger-scale collective motion. To provide a microscopic explanation of this behavior, we present the location of the topological defects (TDs) in the same panels, using red symbols for vortices and blue symbols for anti-vortices. We notice that at short times ($\tau=0.05$ s), the topological defects are located in random positions, forming a gas of defects. Moreover, TDs tend to be located in regions with low polarization, indicating that TDs are microscopic objects that disrupt the onset of collective motion by randomizing the direction of the particle displacements. More details on the spatial correlation between TDs as a function of the observation time $\tau$ can be found in the Supplementary Information.
@@
Importantly, we observe that the onset of collective motion, characterized by the percolation of clusters with large $P_{c}$, is mediated by the annihilation of topological defects that ‘live’ at the interfaces between these collective clusters (blue regions). This annihilation occurs between pairs of defects with opposite charges, preserving the local winding number. Two examples of this process are shown with enlarged views in Fig.3(b). It is also evident and discussed in more detail below that the number of TDs decreases with the observation time, favoring the establishment of large-scale collective clusters.
Collective motion provides another manifestation of dynamical heterogeneity. In Fig. S4(b) of the Supplementary Information, we compute the four-point susceptibility $\chi_{4}(t)$ to track its evolution over the observation time. The absence of a peak in $\chi_{4}(t)$ within the experimental run suggests that the characteristic timescale of these heterogeneities is at least several thousand seconds—much longer than the observation window considered in this study.
Previous work by Silke Henkes and colleagues also demonstrated large-scale correlated dynamics reminiscent of dynamical heterogeneities [78, 79]. However, and crucially different from our case, their system corresponds to a jammed solid phase, whose dynamics are dominated by elasticity.
@@
To provide a complementary view of these dynamics, in Fig.3(c), we present the same particle displacements and TDs but, this time, superimposed onto the amplitude of the local displacement curl $\Omega_{l}$. As expected, the local curl color map (Fig.3(b)) and the local polarization color map (Fig.3(a)) are roughly complementary. Moreover, the local curl $\Omega_{l}$ is significant around the TDs, that can therefore be identified as the microscopic objects disrupting the irrotational nature of the displacement field. In this complementary view, the onset of collective motion is driven by the percolation of areas with low local displacement curl, which is associated with the annihilation of defect pairs (Fig.3(c)).
@@
Figure 5: Three stage non-equilibrium dynamics at the packing fraction of 0.822. (a): Topological signatures. Defect statistics as a function of the observation time $\tau$. $N_{\text{defects}}$ is the total number of TDs. The vertical dashed lines indicate the two time scales $t_{1}=3\,s$, $t_{2}=60\,s$. (b): Particle-level dynamics. Mean square displacement (MSD) in units of the square of the small particle diameter $D_{s}^{2}$. (c): Collective motion. Average curl of the system $\Omega_{0}$. The three red arrows at each panel correspond to the selected times shown in Fig.3, while the colored backgrounds indicate the system’s dynamics across the three stages.